OpenStudy (anonymous):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal to the nearest tenth? Data Set 1: {85, 83, 90, 70, 87} Data Set 2: {75, 74, 68, 83, 80}

OpenStudy (anonymous):

@Michele_Laino @misty1212

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@radar @thomaster

OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP

OpenStudy (michele_laino):

please wait a moment

OpenStudy (anonymous):

okay :)

OpenStudy (michele_laino):

we have to find the mean values of both data sets, first

OpenStudy (anonymous):

the first one is 83

OpenStudy (michele_laino):

ok! and what is the mean of the second data set?

OpenStudy (anonymous):

the second one is 76

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, we have to compute the MAD of each data set

OpenStudy (michele_laino):

for first data set, we have to compute these quantities: \[\begin{gathered} 85 - 83 = ...? \hfill \\ 83 - 83 = ...? \hfill \\ 90 - 83 = ...? \hfill \\ 70 - 83 = ...? \hfill \\ 87 - 83 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

85−83=2 83−83=0 90−83=7 70−83= -13 87−83=4

OpenStudy (michele_laino):

ok! Now we have to consider the absolute values of those quantities, so the MAD for first data set is given by the subsequent computation: \[\Large MAD = \frac{{2 + 0 + 7 + 13 + 4}}{5} = ...?\] please note that I have considered +13, and not -13, since we have to keep the absolute values

OpenStudy (anonymous):

26/5

OpenStudy (michele_laino):

ok! MAD=26/5=5.2

OpenStudy (michele_laino):

then the ratio mean-to.-MAD, for first data set, is: 83/5.2=...?

OpenStudy (anonymous):

15.96153846

OpenStudy (michele_laino):

ok! 15.96

OpenStudy (michele_laino):

now, we have to compute these quantities: \[\begin{gathered} 75 - 76 = ...? \hfill \\ 74 - 76 = ...? \hfill \\ 68 - 76 = ...? \hfill \\ 83 - 76 = ...? \hfill \\ 80 - 76 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

so is the answer 15.96

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

is it ?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

now we have to compute that above quantities, namely: \[\begin{gathered} 75 - 76 = ...? \hfill \\ 74 - 76 = ...? \hfill \\ 68 - 76 = ...? \hfill \\ 83 - 76 = ...? \hfill \\ 80 - 76 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

75−76=1 74−76=2 68−76=8 83−76=7 80−76=4

OpenStudy (michele_laino):

we have: -1, -2, -8, 7, 4

OpenStudy (michele_laino):

now the MAD for second data set is: \[MAD = \frac{{1 + 2 + 8 + 7 + 4}}{5} = ...?\]

OpenStudy (anonymous):

4.4

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so, the ratio mean-to-MAD for second distribution is: 76/4.4=...?

OpenStudy (anonymous):

17.2

OpenStudy (michele_laino):

better is 17.3 furthermore, the ratio mean-to-MAD for first data set is 16.0, since we have to round off 15.96 to the nearest tenth

OpenStudy (anonymous):

16.0

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

can you help me i answered the questions i just need to see if they are right its 5 questions can you check them

OpenStudy (michele_laino):

ok!